OpenStudy (anonymous):

Solve the following system of equations using elimination: 3x - 4y = 8 10x + 5y = 45 A. (-4, -1) B. (3, 3) C. (4, 1) D. (8, 4)

OpenStudy (zzr0ck3r):

multiply the top equation by -10 and the bottom equation by 3

OpenStudy (anonymous):

Where do I do that?

OpenStudy (anonymous):

You'd do that so that the x's cancel out when you then add the equations together.

OpenStudy (zzr0ck3r):

multiply the top equation by -10 -10(3x-4y=8) = -30x+40y=-80 multiply the bottom equation by 3 3(10x+5y=45)=30x+15y=135 now we have -30x+40y=-80 30x+15y=135 add them 55y=55 y=1 now we know the answer

OpenStudy (anonymous):

C?

OpenStudy (nincompoop):

well let us see if c makes sense. plug the values of c in place of x and y and see if the equations stand or will return a true value.

OpenStudy (anonymous):

Idk maybe D seems better.

OpenStudy (nincompoop):

well if you're going to use that technique, you might as well start from A

OpenStudy (nincompoop):

do you know what you're doing here?

OpenStudy (nincompoop):

did you read and analyze zzrock's reply? it seems to me that you didn't fully comprehend it

OpenStudy (anonymous):

It's C.

OpenStudy (anonymous):

5[3x-4y=8] 4[10x+5y=45] 15x-20y=40 40x-20y=180 --------------- 55x=220 x=4 Plug it back in, substituting it for x. 10(4)+5y=45 40+5y=45 5y=5 y=1

OpenStudy (zzr0ck3r):

the fact that i showed y=1 and there is one choice that meets this criteria should have told you the answer, if you do not understand something you should ask about that something, dont just say you dont get it....and redoing the problem im sure is not going to help if she didnt get what I did.

OpenStudy (anonymous):

You want to make either x or y equal, but opposite signs so you can 'eliminate' them. For instance for your problem, the easiest thing to do was to make both y's 20 since 4 and 5 have a common factor. However, to eliminate them, one has to have the opposite sign. Since there was already a '-4' there's no need to make it negative. Also zzrock; mind yourself. I believe I gave her a more thorough explanation. Don't call me out miss.

OpenStudy (anonymous):

You guys calm down its just a question I know I am not good at math I am sorry. But no need to get mad.

OpenStudy (zzr0ck3r):

no one is mad, at least I am the furthest thing from mad. What I am is someone who wants you too understand the solution. so please ask questions. Its not about people repeating the same thing over and over, its about what @nincompoop said. you need to analyze my solution or any of the other solutions given, and ask a question when you dont know what something means.